OpenStudy (anonymous):

Are the lines for the pair of equations parallel, perpendicular, or neither. y = –2/3x + 1 2x – 3y = –3

OpenStudy (anonymous):

the slope of the first line is \(-\frac{2}{3}\) to find the slope of the second line write \[2x-3y=-3\] \[-3y=-2x-3\] \[y=\frac{2}{3}x+1\] so it has slope \(\frac{2}{3}\) the slopes are not the same, so they are not parallel since \(\frac{2}{3}|\) is not the negative reciprocal of \(-\frac{2}{3}\) they are not perpendicular

OpenStudy (anonymous):

Thanks! So if i had a diffrent problem and the slopes were the same it would be parallel?